40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Feb 19 13:28:44 2010, Koji, Update, General, Prep for Power Supply Stop  screen_shot.png
    Reply  Sun Feb 21 17:44:35 2010, rana, Update, General, Prep for Power Supply Stop  
       Reply  Mon Feb 22 07:25:58 2010, rana, Update, General, Prep for Power Supply Stop  
          Reply  Mon Feb 22 09:45:34 2010, josephb, Update, General, Prep for Power Supply Stop  
             Reply  Mon Feb 22 11:42:48 2010, Koji, Update, General, Prep for Power Supply Stop  
Message ID: 2617     Entry time: Fri Feb 19 13:28:44 2010     Reply to this: 2620
Author: Koji 
Type: Update 
Category: General 
Subject: Prep for Power Supply Stop  

- ETMX/ETMY oplev paths renewed. The nominal gain for ETMY YAW was reversed as a steering mirror has been put.
- Oplevs/QPDs cenrtered except for the MCT QPD.
- SUS snapshots updated
- QPD/Aligment screenshots taken

40m Wiki: Preparation for power supply stop

Attachment 1: screen_shot.png  249 kB  | Show | Show all
ELOG V3.1.3-